\begin{thebibliography}{10}

\bibitem{bai2016global}
Xue-Li Bai, Zheng-Hai Huang, and Yong Wang.
\newblock Global uniqueness and solvability for tensor complementarity
  problems.
\newblock {\em Journal of Optimization Theory and Applications}, 170(1):72--84,
  2016.

\bibitem{bailey1969bounds}
Duane~W Bailey and Douglas~E Crabtree.
\newblock Bounds for determinants.
\newblock {\em Linear Algebra and its Applications}, 2(3):303--309, 1969.

\bibitem{cottle2009linear}
RW~Cottle, JS~Pang, and RE~Stone.
\newblock {\em The linear complementarity problem}.
\newblock SIAM, 2009.

\bibitem{cvetkovic2009new}
Ljiljana Cvetkovi{\'c}, Vladimir Kosti{\'c}, and Sonja Rau{\v{s}}ki.
\newblock A new subclass of h-matrices.
\newblock {\em Applied mathematics and computation}, 208(1):206--210, 2009.

\bibitem{das2016properties}
AK~Das.
\newblock Properties of some matrix classes based on principal pivot transform.
\newblock {\em Annals of Operations Research}, 243(1):375--382, 2016.

\bibitem{das2016generalized}
AK~Das, R~Jana, and Deepmala.
\newblock On generalized positive subdefinite matrices and interior point
  algorithm.
\newblock In {\em International Conference on Frontiers in Optimization: Theory
  and Applications}, pages 3--16. Springer, 2016.

\bibitem{das2017finiteness}
AK~Das, R~Jana, and Deepmala.
\newblock Finiteness of criss-cross method in complementarity problem.
\newblock In {\em International Conference on Mathematics and Computing}, pages
  170--180. Springer, 2017.

\bibitem{das2018invex}
AK~Das, R~Jana, and Deepmala.
\newblock Invex programming problems with equality and inequality constraints.
\newblock {\em Transactions of A. Razmadze Mathematical Institute},
  172(3):361--371, 2018.

\bibitem{das2019some}
AK~Das, R~Jana, and Deepmala.
\newblock Some aspects on solving transportation problem.
\newblock {\em Yugoslav Journal of Operations Research}, 30(1):45--57, 2019.

\bibitem{ding2018p}
Weiyang Ding, Ziyan Luo, and Liqun Qi.
\newblock P-tensors, p0-tensors, and their applications.
\newblock {\em Linear Algebra and its Applications}, 555:336--354, 2018.

\bibitem{ding2013m}
Weiyang Ding, Liqun Qi, and Yimin Wei.
\newblock M-tensors and nonsingular m-tensors.
\newblock {\em Linear Algebra and Its Applications}, 439(10):3264--3278, 2013.

\bibitem{dutta2021some}
A~Dutta and AK~Das.
\newblock On some properties of $ k $-type block matrices in the context of
  complementarity problem.
\newblock {\em arXiv preprint arXiv:2109.09549}, 2021.

\bibitem{dutta2022on}
A~Dutta, R~Jana, and AK~Das.
\newblock On column competent matrices and linear complementarity problem.
\newblock In {\em Proceedings of the Seventh International Conference on
  Mathematics and Computing}, pages 615--625. Springer Singapore, 2022.

\bibitem{garcia2014error}
Marta Garc{\'\i}a-Esnaola and Juan~Manuel Pe{\~n}a.
\newblock Error bounds for linear complementarity problems of nekrasov
  matrices.
\newblock {\em Numerical Algorithms}, 67(3):655--667, 2014.

\bibitem{huang2017formulating}
Zheng-Hai Huang and Liqun Qi.
\newblock Formulating an n-person noncooperative game as a tensor
  complementarity problem.
\newblock {\em Computational Optimization and Applications}, 66(3):557--576,
  2017.

\bibitem{huang2019tensor}
Zheng-Hai Huang and Liqun Qi.
\newblock Tensor complementarity problems—part $\text{III}$: applications.
\newblock {\em Journal of Optimization Theory and Applications},
  183(3):771--791, 2019.

\bibitem{jana2019hidden}
R~Jana, AK~Das, and A~Dutta.
\newblock On hidden ${Z}$-matrix and interior point algorithm.
\newblock {\em Opsearch}, 56(4):1108--1116, 2019.

\bibitem{jana2021iterative}
R~Jana, AK~Das, and VN~Mishra.
\newblock Iterative descent method for generalized leontief model.
\newblock {\em Proceedings of the National Academy of Sciences, India Section
  A: Physical Sciences}, 91(2):237--244, 2021.

\bibitem{jana2018processability}
R~Jana, AK~Das, and S~Sinha.
\newblock On processability of $\mbox{L}$emke's algorithm.
\newblock {\em Applications \& Applied Mathematics}, 13(2), 2018.

\bibitem{jana2018semimonotone}
R~Jana, AK~Das, and S~Sinha.
\newblock On semimonotone star matrices and linear complementarity problem.
\newblock {\em arXiv preprint arXiv:1808.00281}, 2018.

\bibitem{jana2021more}
R~Jana, A~Dutta, and AK~Das.
\newblock More on hidden ${Z}$-matrices and linear complementarity problem.
\newblock {\em Linear and Multilinear Algebra}, 69(6):1151--1160, 2021.

\bibitem{kolotilina2015some}
L~Yu Kolotilina.
\newblock Some characterizations of nekrasov and s-nekrasov matrices.
\newblock {\em Journal of Mathematical Sciences}, 207(5):767--775, 2015.

\bibitem{lim2005singular}
Lek-Heng Lim.
\newblock Singular values and eigenvalues of tensors: a variational approach.
\newblock In {\em 1st IEEE International Workshop on Computational Advances in
  Multi-Sensor Adaptive Processing, 2005.}, pages 129--132. IEEE, 2005.

\bibitem{mangasarian1976linear}
Olvi~L Mangasarian.
\newblock Linear complementarity problems solvable by a single linear program.
\newblock {\em Mathematical Programming}, 10(1):263--270, 1976.

\bibitem{mohan2001more}
SR~Mohan, SK~Neogy, and AK~Das.
\newblock More on positive subdefinite matrices and the linear complementarity
  problem.
\newblock {\em Linear Algebra and its Applications}, 338(1-3):275--285, 2001.

\bibitem{mohan2001classes}
SR~Mohan, SK~Neogy, and AK~Das.
\newblock On the classes of fully copositive and fully semimonotone matrices.
\newblock {\em Linear Algebra and its Applications}, 323(1-3):87--97, 2001.

\bibitem{mohan2004note}
SR~Mohan, SK~Neogy, and AK~Das.
\newblock A note on linear complementarity problems and multiple objective
  programming.
\newblock {\em Mathematical programming}, 100(2):339--344, 2004.

\bibitem{mondal2016discounted}
P~Mondal, S~Sinha, SK~Neogy, and AK~Das.
\newblock On discounted $\mbox{ARAT}$ semi-markov games and its complementarity
  formulations.
\newblock {\em International Journal of Game Theory}, 45(3):567--583, 2016.

\bibitem{neogy2016optimization}
SK~Neogy, RB~Bapat, and AK~Das.
\newblock Optimization models with economic and game theoretic applications.
\newblock {\em Annals of Operations Research}, 243(1):1--3, 2016.

\bibitem{neogy2005linear}
SK~Neogy and AK~Das.
\newblock Linear complementarity and two classes of structured stochastic
  games.
\newblock {\em Operations Research with Economic and Industrial Applications:
  Emerging Trends, eds: SR Mohan and SK Neogy, Anamaya Publishers, New Delhi,
  India}, pages 156--180, 2005.

\bibitem{neogy2005almost}
SK~Neogy and AK~Das.
\newblock On almost type classes of matrices with ${Q}$-property.
\newblock {\em Linear and Multilinear Algebra}, 53(4):243--257, 2005.

\bibitem{neogy2005principal}
SK~Neogy and AK~Das.
\newblock Principal pivot transforms of some classes of matrices.
\newblock {\em Linear algebra and its applications}, 400:243--252, 2005.

\bibitem{neogy2006some}
SK~Neogy and AK~Das.
\newblock Some properties of generalized positive subdefinite matrices.
\newblock {\em SIAM journal on matrix analysis and applications},
  27(4):988--995, 2006.

\bibitem{neogy2008mathematical}
SK~Neogy and AK~Das.
\newblock {\em Mathematical programming and game theory for decision making},
  volume~1.
\newblock World Scientific, 2008.

\bibitem{neogy2011generalized}
SK~Neogy and AK~Das.
\newblock Generalized monotone maps and complementarity problems.
\newblock In {\em Topics in Nonconvex Optimization}, pages 27--46. Springer,
  2011.

\bibitem{neogy2011singular}
SK~Neogy and AK~Das.
\newblock On singular ${N_0}$-matrices and the class ${Q}$.
\newblock {\em Linear algebra and its applications}, 434(3):813--819, 2011.

\bibitem{neogy2013weak}
SK~Neogy and AK~Das.
\newblock On weak generalized positive subdefinite matrices and the linear
  complementarity problem.
\newblock {\em Linear and Multilinear Algebra}, 61(7):945--953, 2013.

\bibitem{neogy2009modeling}
SK~Neogy, AK~Das, and RB~Bapat.
\newblock {\em Modeling, computation and optimization}, volume~6.
\newblock World Scientific, 2009.

\bibitem{neogy2012generalized}
SK~Neogy, AK~Das, and A~Gupta.
\newblock Generalized principal pivot transforms, complementarity theory and
  their applications in stochastic games.
\newblock {\em Optimization Letters}, 6(2):339--356, 2012.

\bibitem{neogy2008mixture}
SK~Neogy, AK~Das, S~Sinha, and A~Gupta.
\newblock On a mixture class of stochastic game with ordered field property.
\newblock In {\em Mathematical programming and game theory for decision
  making}, pages 451--477. World Scientific, 2008.

\bibitem{orera2019accurate}
H~Orera and JM~Pe{\~n}a.
\newblock Accurate inverses of nekrasov z-matrices.
\newblock {\em Linear Algebra and its Applications}, 574:46--59, 2019.

\bibitem{qi2005eigenvalues}
Liqun Qi.
\newblock Eigenvalues of a real supersymmetric tensor.
\newblock {\em Journal of Symbolic Computation}, 40(6):1302--1324, 2005.

\bibitem{shao2013general}
Jia-Yu Shao.
\newblock A general product of tensors with applications.
\newblock {\em Linear Algebra and its applications}, 439(8):2350--2366, 2013.

\bibitem{shao2016some}
Jiayu Shao and Lihua You.
\newblock On some properties of three different types of triangular blocked
  tensors.
\newblock {\em Linear Algebra and its Applications}, 511:110--140, 2016.

\bibitem{song2014properties}
Yisheng Song and Liqun Qi.
\newblock Properties of tensor complementarity problem and some classes of
  structured tensors.
\newblock {\em arXiv preprint arXiv:1412.0113}, 2014.

\bibitem{song2015properties}
Yisheng Song and Liqun Qi.
\newblock Properties of some classes of structured tensors.
\newblock {\em Journal of Optimization Theory and Applications},
  165(3):854--873, 2015.

\bibitem{song2017properties}
Yisheng Song and Liqun Qi.
\newblock Properties of tensor complementarity problem and some classes of
  structured tensors.
\newblock {\em Annals of Applied Mathematics}, 2017.

\bibitem{song2016properties}
Yisheng Song and Gaohang Yu.
\newblock Properties of solution set of tensor complementarity problem.
\newblock {\em Journal of Optimization Theory and Applications}, 170(1):85--96,
  2016.

\bibitem{yuan2014some}
Pingzhi Yuan and Lihua You.
\newblock Some remarks on p, p0, b and b0 tensors.
\newblock {\em Linear Algebra and its Applications}, 459:511--521, 2014.

\bibitem{zhang2018nekrasov}
Junli Zhang and Changjiang Bu.
\newblock Nekrasov tensors and nonsingular $\mathcal{H}$- tensors.
\newblock {\em Computational and Applied Mathematics}, 37(4):4917--4930, 2018.

\bibitem{zhang2014m}
Liping Zhang, Liqun Qi, and Guanglu Zhou.
\newblock M-tensors and some applications.
\newblock {\em SIAM Journal on Matrix Analysis and Applications},
  35(2):437--452, 2014.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath}
\usepackage{float}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage[utf8x]{inputenc}
\usepackage{caption}
\usepackage{graphics}
\usepackage{enumerate} 
\usepackage{algorithmic}
\usepackage{calrsfs}
\DeclareMathAlphabet{\pazocal}{OMS}{zplm}{m}{n}

\usepackage[Algorithm,ruled]{algorithm}
\usepackage{titlesec}
\usepackage{sectsty}
\titleformat*{\section}{\LARGE\bfseries}
\titleformat*{\subsection}{\Large\bfseries}
\titleformat*{\subsubsection}{\large\bfseries}
\sectionfont{\Huge}
\subsectionfont{\large}
\subsubsectionfont{\normalsize}
\usepackage{multicol}
\usepackage{lmodern}
\usepackage{marvosym} 
\usepackage{lipsum}
\usepackage{mwe}
\usepackage{caption}
\usepackage{subcaption} 
\newtheoremstyle{case}{}{}{}{}{}{:}{ }{}
\theoremstyle{case}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\ben}{\begin{eqnarray*}}
\newcommand{\een}{\end{eqnarray*}}
\newtheorem{examp}{\sc example}
\newtheorem{remk}{\sc remark}
\newtheorem{corol}{\sc corollary}
\newtheorem{lemma}{\sc lemma}
\newtheorem{theorem}{\sc theorem}
\newtheorem{defn}{\sc definition}
\newtheorem{prop}{\sc proposition}
\newcommand{\bt}{\begin{theorem}}
\newcommand{\et}{\end{theorem}}
\newcommand{\bl}{\begin{lemma}}
\newcommand{\el}{\end{lemma}}
\newcommand{\bed}{\begin{defn}}
\newcommand{\eed}{\end{defn}}
\newcommand{\brem}{\begin{remk}}
\newcommand{\erem}{\end{remk}}
\newcommand{\bex}{\begin{examp}}
\newcommand{\eex}{\end{examp}}
\newcommand{\bcl}{\begin{corol}}
\newcommand{\ecl}{\end{corol}}

\newcommand{\NI}{\noindent}

\theoremstyle{definition}
\theoremstyle{remark}

\numberwithin{equation}{section}
\numberwithin{theorem}{section}
\numberwithin{lemma}{section}

\begin{document}

\title{\large\bf\sc Complementarity Problem With Nekrasov $Z$ tensor}

\author{ 
R. Deb$^{a,1}$ and A. K. Das$^{b,2}$\\
\emph{\small $^{a}$Jadavpur University, Kolkata, India.}\\	
\emph{\small $^{b}$Indian Statistical Institute, Kolkata, India.}\\
\emph{\small $^{1}$Email: rony.knc.ju@gmail.com}\\
\emph{\small $^{2}$Email: akdas@isical.ac.in}\\
}

\date{}

\maketitle

\begin{abstract}
\NI It is worth knowing that a particular tensor class belongs to $P$-tensor which ensures the compactness to solve tensor complementarity problem (TCP). In this study, we propose a new class of tensor, Nekrasov $Z$ tensor, in the context of the tensor complementarity problem. We show that the class of $P$-tensor contains the class of even ordered Nekrasov $Z$ tensors with positive diagonal elements. In this context, we propose a procedure by which a Nekrasov $Z$ tensor can be transformed into a tensor which is diagonally dominant.\\

\noindent{\bf Keywords:} Diagonally dominant tensor, Nekrasov tensors, Nonsingular $H$ tensor, Nekrasov $Z$ tensor, $P$-tensors, Tensor complementarity problem.\\

\noindent{\bf AMS subject classifications:} 15A69, 99C33, 65F15
\end{abstract}
\footnotetext[1]{Corresponding author}

\section{Introduction}
$\mathcal{T}= (t_{j_1 ... j_m}) $ has multidimensional array of elements $t_{j_1... j_m} \in \mathbb{C}$ where $j_i \in [n]$ with $i\in [m],$ $[n]=\{1,...,n\}.$ The set of $m$th order $n$ dimensional complex (real) tensors is denoted by $\mathbb{C}^{[m,n]}\; (\mathbb{R}^{[m,n]}).$ 
Several classes of tensors have significant impact on different branches of science and engineering. Applications of tensors can be found in physics, quantum computing, diffusion tensor imaging, image authenticity verification problem, spectral hypergraph theory, optimization theory and in several other areas. In optimization theory, Song and Qi \cite{song2017properties} introduced the tensor complementarity problem (TCP) which is identified a subclass of the nonlinear complementarity problem. Here functions are obtained with the help of a tensor.

\NI Given $\mathcal{T}\in \mathbb{R}^{[m,n]}$ and $q\in\mathbb{R}^n,$ TCP is to obtain $x\in \mathbb{R}^n $ such that
\begin{equation}\label{tensor comp equation}
	x\geq 0,\;\;  \mathcal{T}x^{m-1}+q \geq 0,\;\;  x^T (\mathcal{T}x^{m-1}+q)  = 0.
\end{equation}
This problem is denoted by TCP$(\mathcal{T}, q)$ and the solution set of TCP$(\mathcal{T}, q)$ is denoted by SOL$(\mathcal{T}, q).$
TCP are observed in optimization theory, game theory and in other areas. Some mathematical problem can be reformulated TCP, such as a class of multiperson noncooperative game \cite{huang2017formulating}, hypergraph clustering problem and traffic equilibrium problem \cite{huang2019tensor}.

\NI TCP can be viewed as one kind of extension of the linear complementarity problem (LCP) where the involved functions are homogeneous polynomials constructed with the help of a tensor.
\NI Given $M\in \mathbb{R}^{[2,n]}$ and $q\in \mathbb{R}^n$, the LCP \cite{cottle2009linear} is to obtain $u\in \mathbb{R}^n$ such that
\begin{equation}\label{linear comp equation}
 u\geq 0,\;\;  Mu + q \geq 0,\;\; u^T (Mu + q) = 0.
\end{equation}
The problem is denoted by LCP$(M,q)$ and the solution set of LCP$(M,q)$ is denoted by SOL$(M,q).$ The stability of a large number of problems can be explained with the formulation of a linear complementarity problem. For details see \cite{das2019some}, \cite{neogy2006some}, \cite{dutta2021some} \cite{dutta2022on}, \cite{neogy2013weak}, \cite{jana2018semimonotone}, \cite{neogy2005almost}, \cite{neogy2011singular}, \cite{jana2019hidden}, \cite{jana2021more}, \cite{neogy2009modeling}, \cite{das2017finiteness}, \cite{das2018invex}. For the formulation of conflicting situations see \cite{mondal2016discounted}, \cite{neogy2008mathematical}, \cite{neogy2008mixture}, \cite{neogy2005linear}, \cite{neogy2016optimization}, \cite{das2016generalized}, \cite{neogy2011generalized} and  \cite{mohan2004note}. In connection with linear complementarity problem computation methods are dependent on the matrix classes. For details see \cite{mohan2001classes}, \cite{mohan2001more} \cite{neogy2005principal}, \cite{das2016properties}, \cite{neogy2012generalized}, \cite{jana2019hidden}, \cite{jana2021iterative}, \cite{jana2021more}, \cite{jana2018processability}.

\NI Nekrasov $Z$-matrix was introduced by Orea and Pe$\tilde{n}$a \cite{orera2019accurate} to study the methods to compute its inverse with high relative accuracy.

\NI Several matrix types play a crusial role in the theory of linear complementarity problem, similarly, some structured tensors are important to study tensor complementarity theory. In recent years some special types of matrices are extended to tensors in connection with TCP. Among the several tensor classes, the class of $P(P_0)$-tensor, $Z$ tensor, $M$-tensor, $H$ tensor, and Nekrasov tensor are of special interest. $Z$ tensor and $M$-tensor are introduced by Zhang et al. \cite{zhang2014m} in context of tensor complementarity problem. They prove that for a $Z$ tensor the TCP has the least element property. Ding et al. \cite{ding2013m} introduce $H$ tensor and establish its importance in the tensor complementarity problem. Zhang and Bu \cite{zhang2018nekrasov} introduce the Nekrasov tensor and show that a Nekrasov tensor can be posed as a $H$ tensor.

\NI Song and Qi \cite{song2014properties} introduce $P$-tensor. Various properties of $P$-tensors are considered in connection with TCP. For details see \cite{bai2016global}, \cite{ding2018p}, \cite{yuan2014some}.

\NI In this paper, we introduce Nekrasov $Z$ tensor. We study the connection of Nekrasov $Z$ tensor and diagonally dominant tensor. We prove that the class of Nekrasov $Z$ tensor of even order with positive diagonal elements is a subclass of $P$-tensor.

The paper is organized as follows. Section 2 contains few essential and early observations. In Section 3, we develop Nekrasov $Z$ tensor.  We investigate its connection with the $P$-tensor.

\section{Preliminaries}
Here we consider some basic notations and present few useful definitions and theorems used in this paper. $x\in \mathbb{C}^n$ denotes a column vector and row transpose of $x$ is denoted by $x^T.$ A diagonal matrix $W=[w_{ij}]_{n \times n}=diag(w_1, \; ..., \; w_n)$ is defined as $w_{ij}=\left \{ \begin{array}{ll}
	  w_i  &;\; \forall \; i=j, \\
	  0  &; \; \forall \; i \neq j.
	   \end{array}  \right.$

\begin{defn}\cite{mangasarian1976linear}
$A\in \mathbb{R}^{[2,n]},$ is a $Z$-matrix if all its off-diagonal elements are nonpositive.
\end{defn}	   

\begin{defn}\cite{cvetkovic2009new,bailey1969bounds,kolotilina2015some, garcia2014error}
A matrix $M=(m_{ij})\in \mathbb{C}^{[2,n]},$ is a Nekrasov matrix if $\lvert m_{ii} \rvert> \Lambda_i(M), \; \forall\; i\in [n],$ where
\begin{center}
    $\Lambda_i(M) =\left\{\begin{array}{ll}
        \sum_{j=2}^n \lvert m_{ij} \rvert &, i=1,  \\
        \sum_{j=1}^{i-1} \lvert m_{ij} \rvert \frac{\Lambda_j(M)}{\lvert m_{jj} \rvert} + \sum_{j=i+1}^n \lvert m_{ij} \rvert &, i=2,3,...,n.
    \end{array} \right.$
\end{center}
\end{defn}

\begin{defn}\cite{orera2019accurate}
$A\in \mathbb{R}^{[2,n]}$ is Nekrasov $Z$-matrix, if $A$ is a Nekrasov mtrix as well as a $Z$-matrix.
\end{defn}


\noindent Let $\mathcal{O}$ be the zero tensor with zero at each entry of $\mathcal{O}.$ For $\mathcal{T}\in \mathbb{C}^{[m,n]}$ and $x\in \mathbb{C}^n,\; \mathcal{T}x^{m-1}\in \mathbb{C}^n $ is a vector defined by
\[ (\mathcal{T}x^{m-1})_i = \sum_{i_2, ..., i_m =1}^{n} t_{i_1  ...i_m} x_{i_2}  \cdots x_{i_m} , \mbox{   for all } i \in [n] \]
$\mathcal{T}x^m\in \mathbb{C} $ is a scalar defined by
\[ x^T\mathcal{T}x^{m-1}= \mathcal{T}x^m = \sum_{i_1, ..., i_m =1}^{n} t_{i_1  ...i_m} x_{i_1} \cdots x_{i_m}. \]

\NI Let $\mathcal{U}$ and $\mathcal{V}$ be two $n$ dimensional tensor of order $p \geq 2$ and $r \geq 1,$ respectively. Shao \cite{shao2013general} introduced the general product of $\mathcal{U}$ and $\mathcal{V}.$ Let $\mathcal{U} \cdot \mathcal{V} = \mathcal{T}.$ Then $\mathcal{T}$ is an $n$ dimensional tensor of order $((p-1)(r-1)) + 1$ where the elements of $\mathcal{T}$ are given by
\[t_{j \beta_1 \cdots \beta_{p-1} } =\sum_{j_2, \cdots ,j_p \in [n]} u_{j j_2 \cdots j_p} v_{j_2 \beta_1} \cdots v_{j_p \beta_{p-1}},\] where $j \in [n]$, $\beta_1, \cdots, \beta_{p-1} \in [n]^{r-1}$



\begin{defn}\cite{song2016properties}
Given $\mathcal{T} \in \mathbb{R}^{[m,n]} $ and $q\in \mathbb{R}^n$, $x$ solves TCP$(\mathcal{T},q)$ if there exists a $x$ satisfying the equation (\ref{tensor comp equation}).
\end{defn}

\begin{defn}\cite{song2015properties}
$\mathcal{T}\in \mathbb{R}^{[m,n]} $ is a $P$-tensor, if for any $x\in \mathbb{R}^n \backslash \{0\}$, $\exists$ a $j\in [n]$ such that $x_j \neq 0$ and $x_j (\mathcal{T}x^{m-1})_j >0.$
\end{defn}

\NI The row subtensors are defined in \cite{shao2016some}. Here for the sake of convenience we denote $i$th rowsubtensor of $\mathcal{T}$ by $\mathcal{R}_i(\mathcal{T}).$
\begin{defn}\cite{shao2016some}
For each $i$ the $i$th row subtensor of $\mathcal{T}\in \mathbb{C}^{[m,n]}$ is denoted by $\mathcal{R}_i(\mathcal{T})$ and its elements are given as $(\mathcal{R}_i(\mathcal{T}))_{i_2 ... i_m}=(t_{i i_2... i_m})$, where $i_l\in [n]$ and $2\leq l \leq m.$
\end{defn}

\begin{defn}\cite{zhang2018nekrasov}
Let $\mathcal{T}\in \mathbb{C}^{[m,n]}$ such that $\mathcal{T}=(t_{i_1 i_2 ... i_m})$ and
\begin{equation*}
    R_i(\mathcal{T})= \sum_{(i_2 ... i_m) \neq (i ... i)} \lvert t_{i i_2 ... i_m} \rvert,\; \forall \; i.
\end{equation*}
The tensor $\mathcal{T}$ is diagonally dominant (strict diagonally dominant) if $\lvert t_{i ... i} \rvert \geq (>) R_i(\mathcal{T})$ for all $i\in [n].$
\end{defn}

\begin{defn}\cite{zhang2018nekrasov}
$\mathcal{T}\in \mathbb{C}^{[m,n]}$ is quasidiagonally dominant tensor if there exists a diagonal matrix $W=diag(w_1, w_2, \cdots, w_n)$ such that $\mathcal{T} W$ is strictly diagonally dominant tensor, i.e.,
\begin{equation*}
  \lvert t_{i ... i} \rvert w_i^{m-1} > \lvert t_{i i_2 \cdots i_m}\rvert w_{i_2} \cdots w_{i_m}, \;\;\; \forall\; i\in [n]. 
\end{equation*}
\end{defn}

\NI Qi \cite{qi2005eigenvalues} and Lim \cite{lim2005singular} proposed the idea of eigenvalues and eigenvectors in case of tensors. If a pair $(\lambda, x) \in \mathbb{C}\times (\mathbb{C}^n \backslash \ \{0\})$ satisfies the equation $\mathcal{T} x^{m-1} = \lambda x^{m-1},$ then $\lambda$ is an eigenvalue of $\mathcal{T}$, and $x$ is an eigenvector with respect to $\lambda$ of $\mathcal{T}$, for which $x^{[m−1]} = (x_1^{m−1},... , x_n^{m−1})^T.$ The spectral radius of $\mathcal{T}$ is defined as $\rho(\mathcal{T}) = \max\{|\lambda| \; : \; \lambda \text{ is an eigen value of } \mathcal{T}\}.$

\begin{defn}\cite{ding2013m}
$\mathcal{T}$ is a $Z$ tensor if all its off diagonal elements are nonpositive.
\end{defn}
\begin{defn}\cite{ding2013m}
A $Z$ tensor $\mathcal{T}$ is an $M$ tensor if $\exists$ $\mathcal{B}\; \geq \mathcal{O}$ and $s>0$ such that $\mathcal{T}= s \mathcal{I} - \mathcal{B}$, where $s\geq \rho(\mathcal{B})$. $\mathcal{T}$ is a nonsingular $M$ tensor if $s>\rho(\mathcal{B}).$
\end{defn}

\begin{defn}\cite{ding2013m}
For a tensor $\mathcal{T}= t_{i_1 i_2 ... i_m}\in \mathbb{C}^{[m,n]},$ $\mathcal{C}(\mathcal{T})= c_{i_1 i_2 ... i_m}$ is a comparison tensor of $\mathcal{T}$ if 
\begin{center}
    $c_{i_1 i_2 ... i_m}=\left\{\begin{array}{ll}
       |t_{i_1 i_2 ... i_m}|  & , \text{ if } (i_1, i_2, ... ,i_m) =(i,i, ..., i), \\
       -|t_{i_1 i_2 ... i_m}|  & , \text{ if } (i_1, i_2, ... ,i_m) \neq (i,i, ..., i).
    \end{array} \right.$
\end{center}
\end{defn}

\begin{defn}\cite{ding2013m}
$\mathcal{T}$ is an $H$ tensor, if its comparison tensor is an $M$-tensor, and it is called a nonsingular $H$ tensor if its comparison tensor is a nonsingular $M$-tensor.
\end{defn}

\begin{defn}\cite{zhang2018nekrasov}
For $\mathcal{T}=(t_{i_1 i_2 ... i_m}) \in \mathbb{C}^{[m,n]}, \; t_{i i ... i}\neq 0,$ $\forall \;i$ and
\begin{align*}
    \Lambda_1(\mathcal{T}) & =R_1(\mathcal{T}),\\ 
    \Lambda_i(\mathcal{T}) & =\sum_{i_2...i_m \in [i-1]^{m-1}} \lvert t_{i i_2 ... i_m} \rvert \left(\frac{\Lambda_{i_2} (\mathcal{T})}{\lvert t_{i_2  ... i_2} \rvert}\right)^{\frac{1}{m-1}} \cdots \left(\frac{\Lambda_{i_m} (\mathcal{T})}{\lvert t_{i_m  ... i_m} \rvert}\right)^{\frac{1}{m-1}}\\ 
    & \qquad + \sum_{i_2...i_m \notin [i-1]^{m-1}, (i_2 ... i_m)\neq (i ... i)} \lvert t_{i i_2 ... i_m} \rvert, \;\;\;\; i=2,3,...,n. 
\end{align*}
$\mathcal{T}=(t_{i_1 i_2 ... i_m})\in \mathbb{R}^{[m,n]}$ is a Nekrasov tensor if 
 \begin{equation}
     \lvert t_{i  ... i} \rvert > \Lambda_i(\mathcal{T}), \; \forall\; i\in [n].
 \end{equation}
\end{defn}

\begin{theorem}\cite{ding2013m}\label{nonsingular H if and only if quasi-SDD}
$\mathcal{T}$ is a nonsingular $H$ tensor iff it is a quasi-strictly diagonally dominant tensor.
\end{theorem}

\begin{theorem}\cite{zhang2018nekrasov}\label{nekrasov implies nonsingular H}
If a tensor $\mathcal{T}=(t_{i_1 i_2 ... i_m})\in \mathbb{C}^{[m,n]}$ is a Nekrasov tensor, then $\mathcal{T}$ is a nonsingular $H$ tensor.
\end{theorem}

\begin{theorem}\cite{ding2018p}\label{nonsingular H implies P}
A nonsingular $H$ tensor of even order with all positive diagonal elements is a $P$-tensor.
\end{theorem}

\begin{theorem}\cite{bai2016global}
For $q\in \mathbb{R}^n$ and a $P$-tensor $\mathcal{T} \in \mathbb{R}^{[m,n]}$, the solution set of TCP$(\mathcal{T},q)$ is nonempty and compact.
\end{theorem}

\section{Main results}

We begin by introducing Nekrasov $Z$ tensor.

\begin{defn}
$\mathcal{T}\in \mathbb{R}^{[m,n]}$ is a Nekrasov $Z$ tensor if $\mathcal{T}$ is a Nekrasov tensor as well as a $Z$ tensor.
\end{defn}

The following is an example of a Nekrasov $Z$ tensor.

\begin{examp}\label{example of nekrasov Z tensor}
We consider the Example 3.3 of \cite{zhang2018nekrasov}. Let $\mathcal{B}\in \mathbb{R}^{[4,4]}$ be such that $b_{1111}=8,\; b_{2222}=3.8,\; b_{3333}=3,\; b_{4444}=10,\; b_{1112}=b_{2111}=b_{1211}=b_{1121}=-1,\;$ $b_{3222}=b_{2322}=b_{2232}=b_{2223}=-1,$ $b_{4441}=b_{4414}=b_{4144}=b_{1444}=-3$ and all other elements of $\mathcal{B}$ are zeros. Then $R_1(\mathcal{B})=6,\; R_2(\mathcal{B})=4,\; R_3(\mathcal{B})=1,\; R_4(\mathcal{B})=9.$ Since $|b_{2222}=| < R_2(\mathcal{B}),$ so $\mathcal{B}$ is not diagonally dominant tensor. However $\lambda_1(\mathcal{B})=6,\; \lambda_2(\mathcal{B})=3.75,\;\lambda_3(\mathcal{B})=0.98,\;\lambda_4(\mathcal{B})=9.$ Also $|b_{ii...i}|>\lambda_i(\mathcal{B}),\; \forall\; i\in [4].$ Hence $\mathcal{B}$ is Nekrasov tensor. Also all the off-diagonal elements of $\mathcal{B}$ are nonpositive. Therefore $\mathcal{B}$ is a $Z$ tensor. Hence $\mathcal{B}$ is a Nekrasov $Z$ tensor.
\end{examp}

Now we show that $P$-tensor contains even order Nekrasov $Z$ tensor with positive diagonal elements.

\begin{theorem}\label{1st theorem of Nekrasov Z tensor}
A Nekrasov $Z$ tensor of even order with positive diagonal elements is a $P$-tensor.
\end{theorem}
\begin{proof}
Let $\mathcal{T}$ be a Nekrasov $Z$ tensor which has positive diagonal elements. By Theorem \ref{nekrasov implies nonsingular H}, a Nekrasv tensor is a $H$ tensor which is also nonsingular. Therefore $\mathcal{T}$ is a nonsingular $H$ tensor which has positive diagonal elements. Therefore by Theorem \ref{nonsingular H implies P} $\mathcal{T}$ is a $P$-tensor. 
\end{proof}

Since a Nekrasov tensor is a nonsingular $H$ tensor, then there exists a positive diagonal matrix $W$ such that $\mathcal{T}$ is a quasi-diagonally dominant tensor. Here we are developing a diagonal matrix $W$ for which $\mathcal{T}W$ is a diagonally dominant tensor.

\begin{theorem}
Suppose $\mathcal{T}\in \mathbb{R}^{[m,n]},$ $m$ is even. Let $\mathcal{T}$ be a Nekrasov $Z$ tensor whose diagonal elements are positive and let $W$ be a diagonal matrix such that,
\begin{equation}
    W= diag\left( \left(\frac{\Lambda_1(\mathcal{T})}{t_{11\cdots 1}}\right)^{\frac{1}{m-1}},\;  \left(\frac{\Lambda_2(\mathcal{T})}{t_{22\cdots 2}}\right)^{\frac{1}{m-1}},\; \cdots\; , \left(\frac{\Lambda_n(\mathcal{T})}{t_{nn\cdots n}}\right)^{\frac{1}{m-1}} \right)
\end{equation}
Then $\mathcal{T}W$ is a diagonally dominant $Z$ tensor.
\end{theorem}
\begin{proof}
Firstly note that $\frac{\Lambda_i(\mathcal{T})}{t_{ii\cdots i}} \geq 0$ and so $\left(\frac{\Lambda_i(\mathcal{T})}{t_{ii\cdots i}}\right)^{\frac{1}{m-1}} \geq 0, \; \forall\; i\in [n],$ as $\mathcal{T}$ is a Nekrasov $Z$ tensor with positive diagonal elements and $m$ is even. So $W$ is a nonnegative diagonal matrix. Let $\mathcal{B}=(b_{i_1 i_2 ...i_m})\in \mathbb{R}^{[m,n]}$ be such that $\mathcal{B} = \mathcal{T}W.$ Then
\begin{equation}
    b_{i_1 i_2 ... i_m}= t_{i_1 i_2 ... i_m} \left(\frac{\Lambda_{i_2}(\mathcal{T})}{t_{{i_2}{i_2}\cdots {i_2}}} \right)^{\frac{1}{m-1}} \cdots \left(\frac{\Lambda_{i_m}(\mathcal{T})}{t_{{i_m}{i_m}\cdots {i_m}}} \right)^{\frac{1}{m-1}}
\end{equation}
and
\begin{equation}
    b_{ii ...i} =\Lambda_i(\mathcal{T}).
\end{equation}
It is easy to observe that the signs of $\mathcal{B}$ are same as the signs of $\mathcal{T}.$ This infers that $\mathcal{B}$ is a $Z$ tensor, since $\mathcal{T}$ is a $Z$ tensor. Now we show that $\mathcal{B}$ is a  diagonally dominant tensor. We write
\begin{equation}\label{for the conclusion of Nekrasov Z}
   t_{k\cdots k}= \lvert t_{k\cdots k} \rvert > \Lambda_k(\mathcal{T}), \;\; \forall\; k\in [n],
\end{equation}
since $\mathcal{T}$ is a Nekrasov tensor.
Now for each $i\in [n]$ we obtain,
\begin{align*}
  |b_{i\cdots i}| &=b_{i\cdots i}\\
                  &=\Lambda_i (\mathcal{T})\\
                  &=\sum_{i_2 ... i_m \in [i-1]^{m-1}} \lvert t_{i i_2 ... i_m} \rvert \left(\frac{\Lambda_{i_2} (\mathcal{T})}{\lvert t_{i_2  ... i_2} \rvert}\right)^{\frac{1}{m-1}} \cdots \left(\frac{\Lambda_{i_m} (\mathcal{T})}{\lvert t_{i_m  ... i_m} \rvert}\right)^{\frac{1}{m-1}}\\
                  & \qquad + \sum_{i_2 ... i_m \notin [i-1]^{m-1}, (i_2 ... i_m)\neq (i ... i)} \lvert t_{i i_2 ... i_m} \rvert\\
                  &\geq \sum_{i_2 ... i_m \in [i-1]^{m-1}} \lvert t_{i i_2 ... i_m} \rvert \left(\frac{\Lambda_{i_2} (\mathcal{T})}{\lvert t_{i_2  ... i_2} \rvert}\right)^{\frac{1}{m-1}} \cdots \left(\frac{\Lambda_{i_m} (\mathcal{T})}{\lvert t_{i_m  ... i_m} \rvert}\right)^{\frac{1}{m-1}} \\
                  & \qquad + \sum_{i_2 ... i_m \notin [i-1]^{m-1}, (i_2 ... i_m)\neq (i ... i)} \lvert t_{i i_2 ... i_m} \rvert \left(\frac{\Lambda_{i_2} (\mathcal{T})}{\lvert t_{i_2  ... i_2} \rvert}\right)^{\frac{1}{m-1}} \cdots \left(\frac{\Lambda_{i_m} (\mathcal{T})}{\lvert t_{i_m  ... i_m} \rvert}\right)^{\frac{1}{m-1}} \text{ by (\ref{for the conclusion of Nekrasov Z})} \\
                  &=\sum_{i_2 ... i_m \in [n]^{m-1}, (i_2 ... i_m)\neq (i ... i) } \lvert t_{i i_2 ... i_m} \rvert \left(\frac{\Lambda_{i_2} (\mathcal{T})}{\lvert t_{i_2  ... i_2} \rvert}\right)^{\frac{1}{m-1}} \cdots \left(\frac{\Lambda_{i_m} (\mathcal{T})}{\lvert t_{i_m  ... i_m} \rvert}\right)^{\frac{1}{m-1}}\\
                  &=\sum_{(i_2 ... i_m)\neq (i ... i) } |b_{i i_2 ... i_m}|.
\end{align*}
This implies that $\mathcal{B}$ is a diagonally dominant tensor.
\end{proof}

Now we want to find a class of tensors that will contain Nekrasov $Z$ tensors with positive diagonal elements. To do so we use a decomposition of tensors which will be useful in our discussion.\\

\NI Given a tensor $\mathcal{T}=(t_{i_1 i_2 ... i_m})\in \mathbb{R}^{[m,n]},$ we can write $\mathcal{T}$ as
\begin{equation}\label{decomposition equation}
    \mathcal{T} = \mathcal{B}^+ + \mathcal{C},
\end{equation}
where $\forall \; i\in [n]$ the elements of rowsubtensors of $\mathcal{B}$ and $\mathcal{C}$ are given by,
\begin{equation}\label{law for B+}
(R_i(\mathcal{B}^+))_{i_2 \cdots i_m} = t_{i i_2 ... i_m} - r_i^+, \;\forall \; i_l\in [n], \; l\in [m]
\end{equation}
and
\begin{equation}\label{law for C}
(R_i(\mathcal{C}))_{i_2 \cdots i_m} =r_i^+ , \;\forall \; i_l\in [n], \; l\in [m]..
\end{equation}
with
\begin{equation}\label{law for r_i^+}
    r_i^+ =\max_{i_2, ..., i_m \in [n], \; ( i_2 \cdots i_m) \neq (i \cdots i)} \{ 0, t_{i i_2 ... i_m} \}
\end{equation}

\begin{prop}
Let $\mathcal{T}=(t_{i_1 i_2 ... i_m})\in \mathbb{R}^{[m,n]}.$ If $\mathcal{T}$ is decomposed as given in the equation (\ref{decomposition equation}), then $\mathcal{B}^+$ is a $Z$ tensor and $\mathcal{C}$ is nonnegative tensor.
\end{prop}
\begin{proof}
Firstly, by the definition of the tensor $\mathcal{B}^+$ we have $b_{i_1 i_2 ... i_m} = t_{i_1 i_2 ... i_m} - r_{i_1}^+,$ for all $i_1,i_2, ... ,i_m \in [n].$
Now from the definition of $r_{i_1}^+$ given by equation (\ref{law for r_i^+}) for all $i_1, i_2, ... ,i_m \in [n],$ we write
\begin{align*}
    b_{i_1 i_2 ... i_m} &= t_{i_1 i_2 ... i_m} - r_{i_1}^+\\
    &= t_{i_1 i_2 ... i_m} - \max_{i_2, ..., i_m \in [n], \; ( i_2 \cdots i_m) \neq (i \cdots i)} \{ 0, t_{i i_2 ... i_m} \} \\
    &\leq 0, \text{ when } ( i_2 \cdots i_m) \neq (i \cdots i).
\end{align*}
Therefore all off diagonal elements of the tensor $\mathcal{B}^+$ are nonpositive. Hence the tensor $\mathcal{B}^+$ is a $Z$ tensor.

Secondly, observe that the elements of $\mathcal{C}$ are $r_i^+,$ $\forall \;i\in[n].$ Now by equation (\ref{law for r_i^+}) we write
\begin{equation*}
     r_i^+ =\max_{i_2, ..., i_m \in [n], \; ( i_2 \cdots i_m) \neq (i \cdots i)} \{ 0, t_{i i_2 ... i_m} \}\geq 0.
\end{equation*}
 This implies $\mathcal{C}$ is a nonnegative tensor.
\end{proof}

\begin{remk}
If the tensor $\mathcal{B}^+$ is a Nekrasov tensor then $\mathcal{B}^+$ becomes a Nekrasov $Z$ tensor.
\end{remk}

\begin{remk}
The class of tensors $\mathcal{T}$ that can be expressed as $\mathcal{T} = \mathcal{B}^+ + \mathcal{C},$ where $\mathcal{B}^+$ and $\mathcal{C}$ are constructed by (\ref{law for B+}), (\ref{law for C}) and (\ref{law for r_i^+}) with $\mathcal{B}^+$ as a Nekrasov tensor contains the class of Nekrasov $Z$ tensors.
\end{remk}

\begin{remk}
The $P$ tensors play a crucial role in tensor complementarity theory. We know that the SOL$(\mathcal{T}, q)$ is nonempty and compact if $\mathcal{T}$ is a $P$ tensor. Here we show that $\mathcal{T},$ a Nekrasov $Z$ tensor of even order which has positive diagonal elements, is a $P$-tensor. Hence we conclude that the solution set of TCP$(\mathcal{T}, q)$ is nonempty and compact.
\end{remk}

\section{Conclusion}
In this article, we introduce Nekrasov $Z$ tensor. For a Nekrasov $Z$ tensor $\mathcal{T}$ we find $W$ such that $\mathcal{T}W$ is a diagonally dominant tensor. We prove that the Nekrasov $Z$ tensors of even order with positive diagonal elements are contained in the class of $P$-tensor.

\section{Acknowledgment}
The author R. Deb acknowledges CSIR, India, JRF scheme for financial support.

\bibliographystyle{plain}
\bibliography{referencesall}

\end{document}